# Towards AI Accountability Infrastructure: Gaps and Opportunities in AI Audit Tooling

### v1

@@ 4.4. Performance Analysis
Tools for Performance Analysis ($N=116$) are designed to help auditors evaluate and explain model behavior, usually through the calculation of quantiative metrics related to accuracy/safety ($N=63$), explainability ($N=33$), or fairness ($N=27$). This category includes tools for Fairness Evaluation and Accuracy Evaluation (including tools for A/B testing, benchmarking, and model monitoring, such as Meta’s Dynabench or the Linux Foundation’s Adversarial Robustness Toolbox) as well as Explainability (tools for explaining the behavior of a model, such as IBM’s AI Explainability 360, or for exploring training data, such as Hugging Face’s ROOTS search tool (piktus_roots_2023, 75)) and some tools specifically for Qualitative Analysis.
@@
Despite the many tools developed for these purposes—the most popular AI-specific Github repositories (e.g., OpenAI Evals (openai_gpt-4_2023, 73)) in our survey were for Performance Analysis (Table 4)—practitioners expressed a need for more robust, well-vetted tools and methodologies.
@@
Concerns about methodological integrity. Practitioners had concerns about the validity, reproducibility, transparency, and trustworthiness of the methods used in popular Performance Analysis tools. One auditor at a tech startup said, “I’m still not convinced of the validity, even, of some of those methods” (P4) used in “auditing tools for monitoring and validation. For example, the most popular Performance Analysis tool we found on Github was SHAP (SHapley Additive exPlanations), a game-theoretic method for measuring feature importance in a model (lundberg_unified_2017, 64) (see Table 4). But as (kumar_problems_2020, 55, ) argue, Shapley values are prone to misuse and may be unsuitable for procedural justice, recourse, or normative evaluation. Interpretability and explainability methods promoted by popular tools vary widely in their goals (lipton_mythos_2018, 63) and may encourage false confidence in their users (ghassemi_false_2021, 40), like many “snake oil” AI products (kaltheuner_fake_2021, 50, 70, 97)—yet explainability tools such as SHAP are often suggested in official regulatory guidance (kaye_risky_2023, 52).
@@
Some participants put methodological deficiencies down to differences in the rigor employed by the various disciplines involved with auditing. An auditor in civil society said, “The bar of the kind of threshold of… validity of findings and novelty of findings is much higher in academia than it is for civil society” (P24). Other concerns stemmed from observations about the lack of maintenance of existing tools, the effectiveness of automated monitoring processes, and the efficacy of simulated data for representing real users instead of functioning as “an academic exercise” (P4). As another put it, “you can perturb all the different inputs you want. But they might not be realistic combinations of features for people who are actually using the system” (P25).
@@
Need for inspectable, reproducible methods. To allay methodological concerns, several participants emphasized the importance of open-sourcing data and tools for others in the community to inspect. Some of our participants had reproducibility in mind when designing evaluation procedures in industry: “You want to iterate… but you know that also makes the results less reproducible. And are you being deceptive then, if you [refer in published evaluations] to a model that’s different from the one that people analyze?” (P5) However, in our tool survey, Performance Analysis tools were disproportionately proprietary (30.1%), especially tools for Explainability (42.4%), compared to other tools (22.1% overall; Fig. C.4).
@@
Need for more expansive evaluations. Similar to Standards Identification & Management, our participants wished for tools to help evaluate a broader spectrum of criteria beyond explainability and fairness. Performance Analysis tools—the most popular of which are built by disproportionately disproportionately large ($>10,000$ employees), for-profit firms (Figures C.7)—often focused on a narrow set of technical fairness definitions and explainability methods popularized in academic literature. The tools we surveyed in this category often overlooked entire other areas of concern, including the basic functionality of the model (raji_fallacy_2022, 79) as well as other methods of evaluation. For example, tools specifically devoted to qualitative—as opposed to quantitative—analysis were much harder to find ($N=2$) and rarely mentioned by our participants.
@@
Implications. While there are multiple studies on the use of tools for fairness evaluation (holstein_improving_2019, 48, 58, 30) and explainability (bertrand_selective_2023, 12, 106, 62, 51, 95, 53), fewer studies examine how practitioners evaluate other criteria such as basic functionality (raji_fallacy_2022, 79), safety, privacy, or recourse, just as fewer tools exist for this purpose. Future work should develop and investigate tools for a broader range of evaluation criteria. Also, although our study surfaced many performance analysis tools, practitioners expressed concerns about methodological integrity. Future work should explore practitioners’ standards for audit tooling (kaye_risky_2023, 52), and policymakers may develop standards that require academic peer review or vetting by regulatory bodies for audit tooling.
@@
Moreover, research must examine further how tools may contribute to “audit washing,” the use of auditing procedures to legitimize unethical practices (power_audit_1999, 76). A tool for accuracy evaluation, for example, may be used to analyze the accuracy of dubious technology for predicting “criminality” or “trustworthiness” without questioning underlying ethical issues with these applications (stark_physiognomic_2021, 97). In general, tools may claim to provide auditing capabilities—using terms such as fairness, safety, or explainability—while failing to conduct meaningful evaluations. (wong_seeing_2023, 111, ), for example, show that AI ethics toolkits frame ethics work as primarily technical work for individual practitioners while failing to contend with power dynamics and institutional barriers to accountability.
@@ 4.5. Audit Communication & Advocacy
Some of the most crucial work of an AI audit comes after empirical evaluation is complete. We found two emerging sets of tools that begin to address this important stage of AI auditing: tools for Audit Communication, to effectively translate audit results to a broader audience, and tools for Advocacy, for reporting and campaigning for consequential outcomes in response to audit results. Tools to facilitate Audit Communication were rare in our dataset ($N=2$), all tools for Dataset Visualization (for communicating results in a more accessible way; Google’s FACETS (noauthor_facets_2023, 37), for example). Tools for Advocacy were more common ($N=13$) but still fewer than other stages of our taxonomy, including Community Spaces (e.g., the Benefits Tech Advocacy Hub (noauthor_benefits_2023, 10), which facilitates collaboration and shared resources between advocates who fight algorithm-based cuts to public benefits), tools for Organizing (e.g., the Algorithmic Ecology framework (stop_lapd_spying_coalition_algorithmic_2020, 98), a tool for mapping the non-technical dimensions of algorithmic impact), tools for Legal Search (e.g., the generic case database Westlaw can be used to identify relevant precedent for legal redress), and tools for Resistance (e.g., the gig work app Para, which helps drivers avoid bad algorithmic offers).
@@
These tools were also mentioned less often in our interviews, compared to preceding stages of our taxonomy. Still, auditors were concerned with following through on their evaluation work. As one put it, “in the business of designing audits, it should be as important to design the consequences and penalties that accompany these audits” (P2). Auditors identified many barriers unrelated to tooling that could hinder the impact of audit results, particularly lack of funding. For the auditors we spoke to, tooling in this stage of auditing was mostly aspirational.
@@
Tools and resources for community building. Auditors especially wished for resources that would bring together the diverse, interdisciplinary groups involved in auditing, similar to the few tools for Community Spaces we found in our tool survey. Auditors hoped greater communication could help unite the profession around policy developments, shared language, standards, and goals that could improve the impact of their work. As one auditor put it, “We have to have people in the accountability business” (P7). Another asked, “How do we bring these interdisciplinary communities together so that we can use tools together?” (P20)
Participants also emphasized the importance of establishing guidance and facilitating education about auditing tools. One auditor at a tech startup said, “[NIST] has like AI guidelines that come out, and we work with them… we send in comments, we give talks, all that kind of stuff. I think that’s an important part of the auditing community” (P4). Another described a workshop attended by civil society, academics, and consulting firms to help prepare for legislation in the European Union (P23).
Implications. Communicating audit findings, lessons, and insights learned can help build trust and validate audit findings, recommendations, and subsequent interventions. Published audit reports expand the forum holding model operators accountable, allowing policymakers, journalists, and other public stakeholders to engage with and analyze the outcomes without performing evaluations themselves. Embracing public evaluation results also helps to foster and expedite a culture of continuous learning, enabling audit practitioners to learn from each other’s experiences, and critiques.
Despite these benefits, tooling to support these stages in our taxonomy was rare. While we found some domain-specific spaces where auditors can interact—the Benefits Tech Advocacy Hub (noauthor_benefits_2023, 10), for example—we did not find any audit report registries or other platforms for hosting audit results and facilitating communication between auditors, journalists, and activists. Following calls for toolkits to embrace the non-technical dimensions of AI ethics work (wong_seeing_2023, 111), future design work and research could explore these emergent categories of tooling.
Future audit work should also consider publishing concrete recommendations or demands to facilitate public pressure and accountability, which prior work shows is incredibly impactful in determining the effectiveness of audit target responses (raji_actionable_2022, 78). Demands could include calls for recalls, product updates, or adjustments to communication in marketing materials or terms of service. Academic research could explore in more detail the specific mechanisms of audit communication that are most likely to result in meaningful change, either through voluntary self-correction or through the influence of a regulator or other authority.
5. Overall Observations & Discussion
@@
In this section, we discuss potential paths forward to address the concerns and challenges raised by the current AI audit tool ecosystem. We specifically discuss the important takeaways for the HCI community—which has historically played a key role in the design and development of AI audit tools (sandvig_auditing_2014, 84, 58, 48, 111)—and broader lessons for other stakeholders, including policymakers, audit practitioners, and funders.
@@ 5.1. Moving beyond evaluation, towards accountability
In interviews, accountability was mentioned repeatedly as a major priority for many AI auditors, but is also seemingly one of the most difficult goals to achieve. This aligns with a past survey of AI audit practitioners (costanza-chock_who_2022, 24), in which over 65% felt that “accountability” (defined as a “commitment from auditee to address problems covered by audit within set time”) was an unmet need in their AI auditing work. However, this translation of AI audit results towards accountability outcomes has been under-prioritized for some time. When it comes to the work on AI audit tooling, our tool survey suggests that the focus of audit tool developers has narrowed in on resources for only a subset of the process outlined in Figure 1—we identified six times as many tools in the evaluation stages of our generalized AI audit process as we did tools for harms discovery, audit communication, or advocacy. And understandably so—these are the stages of the audit process that participants expressed as simultaneously being the least resourced and the least understood, requiring contextual awareness and typically under-studied participatory and community engagement methods (see §4.1, §4.5). Here, we identify several key ways HCI researchers and other stakeholders can help expand AI audit tooling beyond evaluation and towards accountability, particularly by studying and developing tools to support other stages of auditing other than evaluation, by continuing to develop tools that facilitate stakeholder participation, and by encouraging and protecting open source tool development.
5.1.1. Implications for HCI
  • Studying and developing tools for harms discovery, audit communication, and advocacy. Most research on AI audit tools focuses on the toolkits used to evaluate AI systems, particularly with respect to fairness and explainability (bellamy_ai_2018, 9, 95, 51, 12, 112, 65, 58, 48). This research is important, but future work should also explore other kinds of tools that could help AI auditors produce effective audit reports. Our tool survey identifies several categories of tools—particularly tools for Incident Reporting, Education/Awareness, Target Identification, Audit Communication, Resistance, Organizing, and Community Spaces—that are also worthy subjects of research. (feffer_ai_2023, 38, ), for example, explore how the AI Incident Database can be used as an educational tool in the classroom. Future research should also explore some of the other challenges practitioners raised, such as toolkits for evaluating basic functionality or privacy, in addition to fairness and explainability, or the process of translating regulatory requirements into specific evaluation criteria.
  • Developing participatory methods for audit work. Participatory processes can be leveraged throughout the AI audit process—when drafting AI principles and conduct standards, when collecting data for AI audits, when defining metrics for performance analysis, and more. Community participation provides the benefit of revealing potential harms, test cases, and mitigation strategies that may be overlooked by auditors and other stakeholders. Directly engaging the community also fosters widespread awareness and efforts towards mobilization and accountability, therefore defining and strengthening what actions constitute meaningful and accountable organizational expectations. Our external audit participants, in particular, embraced collaborative and participatory approaches that involve outside communities in the audit process. Given the broader calls for a participatory turn in AI development  (delgado_participatory_2023, 27, 54, 13), it is no surprise that this is an increasing focus for AI audit practitioners and audit tool developers as well. Recent HCI work on user auditing (deng_understanding_2023, 29), such as the “WeAudit” tool  (devos_toward_2022, 32), exemplifies this shift and highlights the upcoming challenge of designing for a more participatory AI audit process (shen_everyday_2021, 90, 89, 28).
5.1.2. Broader Implications
  • Open sourcing AI audit tools. Some participants were concerned about the efficacy of many tools for Performance Analysis—particularly tools whose methods were not made public. Unlike standards and identification tools, which were mainly open source (Fig. 3), tools for performance analysis were more likely to be proprietary. Making AI audit tools publicly available enables external collaboration as well as third-party validation of the tools’ processes and findings. Open source tools enable outside organizations, researchers, and other stakeholders to access, assess, and benefit from an organization’s internal tools, leading to a collective understanding of current practices and the improvement of otherwise inaccessible methods. By inviting external scrutiny, AI audit tool developers demonstrate a commitment to transparency and accessibility, fostering trust and participation among themselves, AI auditors and other stakeholders. The benefit of knowledge-sharing that occurs when especially internal tools are made available not only leads to tool improvement but enables the replications required to validate the audit results derived from the use of such tools. Policymakers, foundations, and other stakeholders pursuing audit tool development should prioritize open source licensing, and at the least, make their methods public. Policymakers could also consider legislation that requires the publication of not only audit reports and artifacts but also clear explanations of auditors’ methods and tools.
@@ 5.1.2. Broader Implications
  • Protecting auditors and audit tool builders. Multiple participants expressed concerns about disparate power dynamics between auditors and the organization being audited severely limiting the impact of audit results. Interviewees mentioned experiencing audit results being blocked from publication (P1) or unduly restricted in scope (P13) due to interventions from audit targets. Audit target retaliation and censorship was raised as a risk to both internal auditors (P5), who face the threat of firings, social dismissal or professional demotion  (widder_its_2023, 109, 15), and external auditors (P7), who face the threat of legal action under existing privacy and anti-hacking laws (urman_right_2023, 103, 81). For example, the creators of Ad Observer, one of the Data Collection tools surfaced in our survey, were suspended from Facebook because of their tool (edelson_we_2021, 34). In addition to encouraging tool development, policymakers and other influential stakeholders should also take steps to protect auditors. In addition to legal reforms, stakeholders could offer legal funds or services to provide an opportunity for even less-resourced AI auditors experiencing corporate legal retaliation following the publication of an audit report.
5.2. Moving beyond ad hoc toolkits, towards common infrastructure
Throughout our results, auditors highlighted the context-specific, custom-built nature of their auditing tools. As one participant noted, “Many approaches were not necessarily principled. They were quite ad hoc” (P20). While many open source tools existed, auditors often preferred to build their own tooling solutions: “it’s mostly stuff we just doing ourselves because… if we tried to use the existing stuff, it would just complicate that process, I think” (P25). For some, existing tools were inadequate for the complexity and scale of the systems being evaluated: “Most often we try to use open source tools, but that’s very different than a data pipeline that… curates data on millions of [users] every day” (P3). But auditors often also preferred to build their own tools not only because existing tools didn’t fit their needs, but also to give their evaluations more legitimacy. As one internal auditor noted,

### v3

@@ 4.4. Performance Analysis
Tools for Performance Analysis ($N=129$) are designed to help auditors evaluate and explain model behavior, usually through the calculation of quantiative metrics related to accuracy/safety ($N=71$), explainability ($N=36$), or fairness ($N=30$). This category includes tools for Fairness Evaluation, Accuracy Evaluation (including tools for A/B testing, benchmarking, and model monitoring, such as Meta’s Dynabench or the Linux Foundation’s Adversarial Robustness Toolbox), Explainability (tools for explaining the behavior of a model, such as IBM’s AI Explainability 360, or for exploring training data, such as Hugging Face’s ROOTS search tool (piktus_roots_2023)), and Qualitative Analysis.
@@
Concerns about methodological integrity. Despite the many tools developed for Performance Analysis—including the most popular AI-specific Github repositories in our dataset (e.g., OpenAI Evals (openai_gpt-4_2023); see Table E.2)—practitioners expressed a need for more robust, well-vetted tools and methodologies. Internal auditors in particular had concerns about the validity, reproducibility, transparency, and trustworthiness of the methods used in popular Performance Analysis tools. One auditor at a tech startup said, “I’m still not convinced of the validity, even, of some of those methods” (P4) used in tools for monitoring and validation.
@@
For example, the most popular Performance Analysis tool we found on Github is SHAP (SHapley Additive exPlanations), a game-theoretic method for measuring feature importance in a model (lundberg_unified_2017) (see Table E.2). But as kumar_problems_2020 argue, Shapley values are prone to misuse and may be unsuitable for normative evaluation. Interpretability and explainability methods promoted by popular tools vary widely in their goals (lipton_mythos_2018) and, like many “snake oil” AI products (kaltheuner_fake_2021; narayanan_ai_2024; stark_physiognomic_2021), may encourage false confidence in their users (ghassemi_false_2021). Yet explainability tools such as SHAP are often suggested in official regulatory guidance (kaye_risky_2023).
@@
Some participants put methodological deficiencies down to differences in the rigor employed by the various disciplines involved with auditing. An auditor in civil society said, “The bar of the kind of threshold of… validity of findings and novelty of findings is much higher in academia than it is for civil society” (P24). Participants had concerns about maintenance, effectiveness of automated monitoring processes, and the efficacy of synthetic data for representing real users instead of functioning as “an academic exercise” (P4): “You can perturb all the different inputs you want. But they might not be realistic combinations of features for people who are actually using the system” (P25).
@@
Need for inspectable, reproducible methods. To allay methodological concerns, several participants emphasized the importance of open-sourcing tools for others in the community to inspect. Some of our industry participants had reproducibility in mind when designing evaluation procedures: “You want to iterate… but you know that also makes the results less reproducible. And are you being deceptive then, if you [refer in published evaluations] to a model that’s different from the one that people analyze?” (P5) However, the Performance Analysis tools we found (48.1% of which were built by for-profit organizations) were disproportionately not open source compared to other tools in our dataset, especially tools for Explainability (Fig. E.3).
@@
Need for more analysis tools beyond fairness & explainability. Similar to Standards Identification & Management, our participants wished for tools to help evaluate a broader spectrum of criteria for AI systems. The Performance Analysis tools we found—the most popular of which were built and maintained by disproportionately large, for-profit firms (Figures E.6)—often focused on a narrow set of technical fairness definitions and explainability methods popularized in academic literature. The tools we surveyed in this category often overlooked entire other areas of concern, including the basic functionality of the model (raji_fallacy_2022) as well as other methods of evaluation. For example, tools specifically devoted to qualitative—as opposed to quantitative—analysis were much harder to find ($N=3$) and rarely mentioned by our participants.
@@
Implications. While there are multiple studies on the use of tools for fairness evaluation (holstein_improving_2019; lee_landscape_2021; deng_exploring_2022) and explainability (bertrand_selective_2023; wang_designing_2019; liao_questioning_2020; kaur_interpreting_2020; smith-renner_no_2020; kim_help_2023), fewer studies examine how practitioners evaluate other criteria such as basic functionality (raji_fallacy_2022), safety, privacy, or recourse, just as fewer tools exist for this purpose. Future work could develop and investigate tools for a broader range of evaluation criteria. Future work could also explore practitioners’ standards for audit tooling (kaye_risky_2023), and policymakers may develop standards that require academic peer review or vetting by regulatory bodies for audit tooling.
@@
Moreover, research must examine further how tools may contribute to “audit washing,” the use of auditing procedures to legitimize unethical practices (goodman_ai_2022). A tool for accuracy evaluation, for example, may be used to analyze the accuracy of dubious technology for predicting “criminality” or “trustworthiness” without questioning underlying ethical issues with these applications (stark_physiognomic_2021; wang_against_2023). In general, tools may claim to provide auditing capabilities—using terms such as fairness, safety, or explainability—while failing to conduct evaluations that meaningfully contend with power dynamics and institutional barriers to accountability (wong_seeing_2023).
@@ 4.5. Audit Communication & Advocacy
Some of the most crucial accountability work of an AI audit comes after empirical evaluation is complete. We found two emerging sets of tools that begin to address this important stage of AI auditing: tools for Audit Communication, to effectively translate audit results to a broader audience, and tools for Advocacy, for reporting and campaigning for consequential outcomes in response to audit results. Tools to facilitate Audit Communication were the rarest in our dataset ($N=8$), and consist mostly of tools for Dataset Visualization (e.g. Google’s FACETS (noauthor_facets_2023)) and Audit Reporting (e.g., the ACLU’s repository of NYC Local Law 144 hiring bias audit reports (gerchick_tracking_2024)). We found more tools for Advocacy ($N=14$)—including Community Spaces (e.g., the Benefits Tech Advocacy Hub (noauthor_benefits_2023), which facilitates collaboration between advocates who oppose algorithm-based cuts to public benefits), tools for Organizing/Resistance (e.g., the Algorithmic Ecology framework (stop_lapd_spying_coalition_algorithmic_2020), a tool for mapping the non-technical dimensions of algorithmic impact), and tools for Legal Search (e.g., the generic case database Westlaw often used to identify relevant precedent for legal redress)—but still fewer than we found in other stages of our taxonomy.
@@
These types of tools were also mentioned less often in our interviews, compared to preceding stages of our taxonomy. Still, auditors wanted their evaluation work to inform consequential judgments, in line with our definition of an audit (Def. B.2). As one put it, “in the business of designing audits, it should be as important to design the consequences and penalties that accompany these audits” (P2). For the auditors we spoke to, tooling in this stage of auditing was mostly aspirational.
@@
Tools and resources for community building. Auditors especially wished for resources that would bring together the diverse, interdisciplinary groups involved in auditing, similar to the few tools for Community Spaces we found in our tool survey. As one auditor put it, “We have to have people in the accountability business” (P7). Auditors hoped greater communication could help unite the profession around policy developments, shared language, standards, and goals that could improve the impact of their work. One auditor at a tech startup said, “[NIST] has AI guidelines that come out, and we work with them… we send in comments, we give talks, all that kind of stuff. I think that’s an important part of the auditing community” (P4). Another described a workshop attended by civil society, academics, and consulting firms to help prepare for legislation in the European Union (P23). One auditor hoped that communication could lead to shared tooling: “How do we bring these interdisciplinary communities together so that we can use tools together?” (P20)
Implications. Communicating audit findings, lessons, and insights learned can help build trust and validate audit findings, recommendations, and subsequent interventions. Audit report repositories could expand the forum holding model operators accountable, allowing policymakers, journalists, and other public stakeholders to engage with evaluations more easily. Embracing public evaluation results also helps audit practitioners to learn from each other’s experiences. Despite these benefits, tooling to support these stages is rare. While we found some domain-specific spaces where auditors can interact—the Benefits Tech Advocacy Hub (noauthor_benefits_2023), for example—we found few audit reporting tools or tools for facilitating communication between auditors, journalists, and activists. Future design work and research could explore these emergent categories. Academic research could also explore in more detail the specific mechanisms of audit communication that are most likely to result in meaningful change—such as including concrete demands for action (raji_actionable_2022)—and imagine new tools to support those mechanisms.
5. Discussion
@@
The HCI community has historically contributed key research to the design and development of AI audit tools and helped define the concept of AI auditing (sandvig_auditing_2014; lee_landscape_2021; holstein_improving_2019; wong_seeing_2023). In this section, we specifically discuss the important takeaways for that community in particular, as well as broader lessons for other stakeholders, including policymakers, audit practitioners, and funders.
@@ 5.1. Moving beyond evaluation, towards accountability
costanza-chock_who_2022 found that over 65% of surveyed AI audit practitioners felt that “accountability” (defined as a “commitment from auditee to address problems covered by audit within set time”) was a top unmet need in their AI auditing work. This echoes a theme repeated several times throughout our interviews—AI auditors care deeply about accountability but struggled to achieve it.
Despite searching deliberately for non-evaluation tools, we found more than five times as many tools in the evaluation stages of the AI audit process as we did tools for harms discovery, audit communication, or advocacy. Perhaps unsurprisingly, these are also the stages of the audit process that participants described as most requiring contextual awareness and typically under-studied participatory and community engagement methods. Research and development related to these and other practical challenges could bolster practitioners’ accountability efforts. Promising new directions for HCI research and policy include:
  • Studying and developing tools for harms discovery, audit communication, and advocacy. Our tool survey identifies several neglected categories of tools—particularly tools for Incident Reporting, Education/Awareness, Target Identification, and Audit Communication—that are worthy subjects for future HCI research. For instance, promising recent research explores the existing limitations (turri_why_2023) and educational applications (feffer_ai_2023) of incident reporting databases, but little work explores complementary technical infrastructure, such as, for example, the AI inventories often used by journalists (such as Algorithm Tips) and previously required for federal agencies (biden_executive_2023). Likewise, auditors envisioned audit report databases as accountability tools to facilitate the amalgamation and communication of audit findings to key stakeholders, mirroring interventions such as the U.S. Security and Exchange Commission (SEC) EDGAR database for financial accounting audits  (raji_outsider_2022). These gaps in development also present meaningful opportunities for further investment and institutionalization by policy-makers and funders.
  • Validating existing tools in practice. Audits have limited impact if their results are not reliable or meaningfully connected to real world requirements (raji_actionable_2022). Unreliable performance or accuracy analysis tools that fail to meaningfully assess the audit target operate as misleading “rubber stamps” for vendors and lead to “audit washing” (goodman_ai_2022), posing a serious challenge to the legitimacy of audit results. There is a growing opportunity for HCI researchers to explore ways that AI audit practices interact with existing accountability processes such as litigation or regulatory compliance. For instance, several tools surfaced in our survey (e.g., from Holistic.ai, Credo.ai) were explicitly marketed for use for NYC Local Law 144 compliance, but studies of audit practice suggest the produced measures may not be reliable or legally compatible (xiang_legal_2019; groves_auditing_2024; wright_null_2024). Researchers might also investigate the validity and effectiveness of government-sponsored tooling  (kaye_risky_2023) and court-mandated transparency infrastructure (e.g., Facebook Ad Library, built after settlements with civil rights groups (sandberg_doing_2019)).
  • Developing participatory methods for audit work.
    Given the broader calls for a participatory turn in AI development  (delgado_participatory_2023; kulynych_participatory_2020; birhane_power_2022), it is no surprise that participation is an increasing focus for AI audit practitioners and audit tool developers as well. Recent HCI work on user auditing (deng_understanding_2023), such as the “WeAudit” tool  (devos_toward_2022), exemplifies this shift and demonstrates the possibility of designing for a more participatory AI audit process (shen_everyday_2021; shen_publics---loop_2022; deng_understanding_2023). Policymakers can also further emphasize participation as a requirement in audit guidance and invest in tools that support participatory methods.
  • Open & reproducible practices for AI audit tools. Some participants were concerned about the efficacy of many AI audit tools—particularly tools whose methods were not made available for public scrutiny. (Tools we found for Performance Analysis were less likely to be open source; Fig. 3). Making AI audit tools publicly available enables both external collaboration and third-party validation. Open tooling practices may also contribute to knowledge-sharing, standards-setting, transparency, accessibility, and trust, but can have complex interactions with power and oversight that are worthy of further study (widder_why_2024). Researchers, policymakers, foundations, and other stakeholders developing audit tools should prioritize open practices. Policymakers could also consider requiring that published audit reports include clear explanations of auditors’ methods and tools.
@@
  • Independence in audit tool use and protection from retaliation. Power dynamics between auditors and the audited have a critical impact on accountability (raji_outsider_2022; birhane_ai_2024). Participants had audit results blocked from publication (P1) or unduly restricted in scope (P13) due to interventions from audit targets. Audit target retaliation and censorship was raised as a risk to both internal auditors (P5), who face the threat of firings, social dismissal or professional demotion  (widder_its_2023; boag_tech_2022), and external auditors (P7), who face the threat of legal action under existing privacy and anti-hacking laws (urman_right_2024; raji_outsider_2022). HCI research has explored software engineers’ attempts to act on ethics concerns in the face of similar risks (widder_its_2023; tahaei_privacy_2021); further work could explore auditors’ experiences specifically. Policymakers and stakeholders can take steps to provide protections for auditors through legal reforms (longpre_safe_2024) or legal funds such as the Coalition for Independent Technology Research.
5.2. Moving beyond ad hoc toolkits, towards shared infrastructure
Participants agreed on the need for shared infrastructure that supports the auditing process. As one participant said, tools are “a superpower for journalists, and something that really is the future of accountability… There really needs to be some sort of public infrastructure [for auditing]” (P7). But developing high quality audit tools—even when adapting existing open source tools—took resources, and participants noted a lack of long-term investment: “We need more funding for this space… especially when it comes to infrastructure” (P20).
Currently, much of the funding for even open source audit tools comes from private, for-profit organizations. In our landscape analysis, we found that even the free tools currently dominating the audit tooling landscape were often built by large, for-profit tech companies (Fig. 2)—for example, 9 of the 12 Transparency Infrastructure tools we found were built by for-profit organizations. Audit tools like these can hold great power over the audit process: